How many C atoms are present in the sample of C3H8 with a mass of 6.21g?
1
Expert's answer
2019-09-09T04:53:11-0400

There are 3 atoms of Carbon in 1 molecule of C3H8. So there is three times more of Carbon atoms in 1 mole of C3H8 (6.022*1023 molecules). Find the amount of C3H8 given:

"\nu = m (C_3H_8)/ M(C_3H_8)"

"\nu = 6.21/44.1 = 0.14 (mol)"

Find the number of C atoms:

N = 0.14 * 6.022 * 1023 * 3 = 2.53 * 1023 (u.)

Answer: 2.53 * 1023 atoms are present in the sample of C3H8 with a mass of 6.21g.